y=sin3t find dy/dt


v=11sin4t - 5cos5t, find dv/dt



differentiate with respect to x


y=12e5x


y=ln3x


1
Expert's answer
2021-04-15T07:32:46-0400

Solution.

1)

"y=\sin{3t},\newline
\frac{dy}{dt}=3\cos{3t}."

2)

"v=11\sin{4t}-5\cos{5t},\newline
\frac{dv}{dt}=44\cos{4t}+25\sin{5t}."

3)

"y=12e^{5x},\newline
\frac{dy}{dx}=60e^{5x}."

4)

"y=\ln{3x},\newline
\frac{dy}{dx}=\frac{3}{3x}=\frac{1}{x}."